OpenStudy (anonymous):

How do we calculate the sum of the individual probabilities of disjoint events?

hero (hero):

P(a) + P(b)

OpenStudy (anonymous):

what does the a and b stand for?

hero (hero):

Usually the question would ask what is the probability of a or b happening. Could stand for two different objects of the same kind. For example, suppose you have a fruit basket of 5 apples and 7 oranges. And then you choose 1 piece of fruit. What is the probability that you choose an apple or an orange. Then you would compute the following \[P(\text{a or b}) = \frac{\text{total apples + total oranges}}{\text{total fruit}}\]

hero (hero):

Actually that's a bad example.

OpenStudy (anonymous):

lol

hero (hero):

Okay so maybe you have some pears as well. Like 4 pears

hero (hero):

so then it would be \[\frac{5 + 7}{16}\] or \[\frac{5}{16} + \frac{7}{16}\]

OpenStudy (anonymous):

i think i get it, like the two combinations the 2 fruits and the total of the 2 on the bottom ?

hero (hero):

or \[\frac{3}{4}\]

OpenStudy (anonymous):

but what do the a and b stand for?

OpenStudy (anonymous):

is it like the 2 fruits?

hero (hero):

They stand for subsets of an entire set of objects.

hero (hero):

All the fruit in the basket represents the entire set. a represents all the apples in the set. b represents all the oranges in the set.

OpenStudy (anonymous):

oh,

hero (hero):

Say that set S represents the total fruit in the basket and a = all the apples b = all the oranges c = all the pears You can represent the set in the following manner: \[s = \brace{a_1, a_2, a_3, a_4, a_5, b_1, b_2, b_3, b_4, b_5, b_6, b_7, c_1, c_2, c_3, c_4}\]

hero (hero):

a = {a1, a2, a3, a4, a5} b = {b1, b2, b3, b4, b5, b6, b7} c = {c1, c2, c3, c4}

OpenStudy (anonymous):

i get it..... lol, thanks!